# Diffusion-Reinforcement Learning Hierarchical Motion Planning in Adversarial Multi-agent Games

### v1

@@ IV-B RL Based Adversarial Behavior Learning
$\rho$ is the initial environment distribution for the start and hideout locations ($s_{0}\sim\rho$ is omitted from the derivation), $\pi^{d}$ is the diffusion global planner, and $\mathbf{w}$ is the generated waypoints to be tracked, $t_{g}$ is the number of waypoints. This equation shows that we are able to decompose our adversarial evading learning problem into a sequential optimization problem, including a waypoint generator and a low-level waypoint tracker, both of which aim to maximize the discounted accumulative reward. However, the unconstrained random exploration of these policies will subject to high variance from the partial observability, therefore we confine our exploration in a more reasonable region described by the diffusion policy distribution in the previous section. We set reward terms to guide the evader to reach the waypoint and avoid detection in Equation (2) where $r_{g}$ is the reward given when reaching the waypoint, $r_{adv}$ is the detection penalty, and $r_{d}$ is the distance penalty to the next waypoint.
@@
TABLE I: Diffusion-RL Benchmarks (mean $\pm$ std)
Domain Prisoner Escape Narco Interdiction
Score $\uparrow$ Detection $\downarrow$ Goal-Reach. $\uparrow$ Score $\uparrow$ Detection $\downarrow$ Goal-Reach. $\uparrow$
@@
A-Star Heuristic 0.721$\pm$0.189 0.344$\pm$0.238 0.980$\pm$0.140 0.635$\pm$0.295 0.392$\pm$0.305 0.800$\pm$0.400
RRT-Star Heuristic 0.731$\pm$0.176 0.321$\pm$0.220 0.940$\pm$0.237 0.655$\pm$0.308 0.375$\pm$0.319 0.830$\pm$0.376
VO-Heuristic 0.738$\pm$0.113 0.323$\pm$0.143 0.980$\pm$0.140 0.635$\pm$0.099 0.426$\pm$0.116 1.000$\pm$0.000
@@
DDPG 0.234$\pm$0.342 0.725$\pm$0.404 0.070$\pm$0.255 0.305$\pm$0.321 0.645$\pm$0.374 0.000$\pm$0.000
SAC 0.779$\pm$0.117 0.058$\pm$0.108 0.130$\pm$0.336 0.486$\pm$0.347 0.434$\pm$0.406 0.010$\pm$0.099
Diffusion Only 0.779$\pm$0.092 0.276$\pm$0.116 1.000$\pm$0.000 0.776$\pm$0.127 0.261$\pm$0.148 1.000$\pm$0.000
@@
Diffusion-RL [ours] 0.899$\pm$0.087 0.116$\pm$0.095 0.960$\pm$0.196 0.886$\pm$0.106 0.123$\pm$0.102 0.940$\pm$0.237
Diffusion-RL-Map [ours] 0.941$\pm$0.079 0.068$\pm$0.080 0.980$\pm$0.140 0.943$\pm$0.069 0.066$\pm$0.080 1.000$\pm$0.000
@@ IV-C Cost Map Construction and Inference
Even though we assume that the cameras and the terrain map information is totally unobservable for the evasive agent, we may know that there are some relations between the detection and these parts on the map. For example, the agent is easier to capture around the camera and more difficult to capture in the low visibility region. Therefore, it is possible for us to spatially show detection risks on a posterior task-oriented cost map associated with the policies we derived as a reverse engineering problem.
Algorithm 4 describes our procedure. We run a certain number of episodes after the hierarchy is trained well (Line 4) and simultaneously record the evader location $\mathbf{X}$ and its distance to the closest searching agent $\mathbf{d}$ (Line 4). Then we select out the evader locations $x_{i}$ with which the associated distance, $d_{i}$, is below a risk threshold $\epsilon$ and add a Gaussian distribution whose mean is $x_{i}$ and standard deviation $\sigma$ (Line 4-4). Next, we perform normalization and output the map.
Importantly, end-users can preview the global paths before execution or revise the cost map with the newly added global information, which adds more flexibility and explainability compared with traditional fully MDP based RL algorithms (Line 4). Finally we can select the global path with lowest cost based on the costmap in the inference (Line 4-4).
@@ 2       $\mathbf{x},\mathbf{d}=\mathbf{Env}(s_{0},\pi^{d},\pi^{l*})$
4             if $d_{i}<\epsilon$ then
@@ V Results & Discussion
In this section, we will first benchmark our diffusion-RL framework against both heuristics and learning based baselines. Then we visualize our constructed costmap which can help explain the results and show more advantages of our method. Then we justify our use of the diffusion model as the path generator. Next we validate our hierarchy in the real robot testbed. Finally we include a result discussion section.
V-A Achieving Long-Term Planning and Short-Term Evasion
We propose to use the following metrics to quantitatively evaluate our method and compare it against the baselines: detection, goal-reaching rate, and score ($\uparrow$ means higher is better and $\downarrow$ means lower is better in Table I). Detection is the average number of steps when the evader is detected for each episode. Goal-reaching rate is the number of episodes in which the evader reach the goal in the time limit divided by the total number of episodes tested. The score metric is a weighted reward term that balances the aforementioned two main objectives in our domains: reaching the goal ($\pm$50) and staying undetected by the opposing blue agents (-1 for each detection). Each metric is normalized by the min-max value of each category and we test the algorithms in 100 episodes with different random seeds from the training.
We compare our diffusion+RL method with learning based and non-learning based baselines. A* and RRT* heuristics are from our previous works [14] where the evader follows the A* or RRT* path and show some pre-scripted adversarial behaviors (e.g. heading to dense forest, changing velocity etc.) when the search team is nearby. Velocity obstacle (VO) method is widely used in the dynamic obstacle avoidance tasks which samples a velocity from the collision free velocity space [54]. From Table I, we find these heuristics can reach the goal at a high rate, but can be detected by the search team much more than our methods which leads to a lower score. The main reason is that the evader will be exposed to the unobservable cameras, which may attract the dynamic search team and the man-crafted evasive heuristic is not optimal for escaping. In comparison our approach, which utilizes the diffusion model as a motion planning module for a reinforcement learning agent, outperforms all non-learning baselines in terms of score and enables the agent to remain mostly undetected by the searching team while reaching the hideout (Table I).
Even though the reinforcement learning (DDPG & SAC) approaches directly optimize the score as their reward function, our methods outperform them on average by 109% on the score metric across both domains. At the same time, we find the goal reaching rate of the DDPG and SAC only methods are very low in prisoner escape domain (7%, 13%) and even lower in narco interdiction domain (0%, 1%). Interestingly, the SAC can outperform the DDPG in each metric we have across the two domains, which meets our expectation that SAC has more exploration ability from its stochastic policy and entropy regularization.
The SAC evader can be detected less than our map-based diffusion-RL algorithm by 19.12%, which is interesting and a red herring. Figure 3 shows a red colorbar for the distance between the evader and the closest agent in the pursuit team, and the diamond purple line indicates the diffusion planned global path. From the trajectory, we can see the SAC algorithm will lead the evader to the edge of the map where there is no camera but not the goal hideout. Instead, our method can loosely follow the diffusion waypoints (purple diamond) and simultaneously deviate a little to avoid the search team when they are close (indicated by magnet and orange dashed box). In addition, our method learns to reduce the evader speed when evading, which lowers the probability of detection, but maintains a higher speed when safer.
Figure 3: Our Diffusion-RL trajectory vs SAC trajectory. Colorbar indicates distance between evader and closest pursuer and purple line indicates the diffusion global path. We find that the our method (left) is able to evade the pursuers by deviating and decreasing the agent speed to reduce detection chance while still being able to reach the goal. On the contrary, the SAC policy is unable to reach the goal and only randomly changes its velocity.
V-B Costmap
@@
Figure 4: Path Planning with the Costmap: The costmap (a) is constructed by correlating the agent’s risk of detection to its location on the map. We show in (b) that the agent is able to succesfully identify where the cameras are. Given this costmap, the agent can select a path that best evades high cost regions (c). Additional obstacles can be added ad-hoc (d) and a new path can be chosen (e). The grey areas indicate untraversable obstacles or danger zones.
@@
As discussed in §IV-C, we build a costmap based on the distance between evader and pursuer which implies detection risks. The costmap serves as an interpretable and memory efficient way to allow our RL agent to plan based on past experiences. Figure 4a shows the costmap, where the high intensity region indicates the evader is close to the search team which may lead to a detection. We also find that the potential detection usually happens around the cameras (see Figure 4b), which is reasonable since the searching team will search the region after the evader is detected by the camera. Therefore, Figure 4c shows the evasive agent can select the best global path which receives lowest cost that can best conceal it from the camera array. At the inference stage, we can adjust the costmap with the new information we received. For example, we can add an undesired danger zone for the evader on the costmap (see Figure 4d) such that it will select another way to reach another hideout which will avoid the danger zone but potentially incurring more detection (see Figure 4e). We ablate our model with and without the costmap and show that the detection from map based path selection is further decreased by a half.
V-C Diffusion Paths Validation
@@
The learned diffusion model should be able to generate diverse global paths that satisfy start and terminal constraints such that it can guide the low-level RL to explore high-valued state spaces. Additionally, the diffusion model should be sample-time efficient to accelerate RL training and inference. Figure 5a-5b compares the diffusion global paths trained on the datasets from RRT* and another commonly used path planner A*. We see the diffusion model encodes a more diverse distribution of the paths leading to the final hideouts when using RRT*. Figure 5c and Table 5d shows the planning time for both the diffusion and RRT* algorithms. The diffusion model takes 85.7% less time generating one trajectory compared with RRT*, which requires a map search. Furthermore, the difference is enlarged when planning more paths since we can draw samples from diffusion model in parallel. We generate the results using a machine with 11th Gen Intel Core i9 processor and GeForce GTX 1660 Ti graphic card.
(a) Diffusion Trajectories (A*)
(b) Diffusion Trajectories (RRT*)
(c) Time Comparison
@@
(d) Time Comparison (mean±$\pm$±std)
Figure 5: The paths from the diffusion model trained on RRT* are more diverse than those trained on A* (5a-5b). Additionally, compared to the traditional RRT* planner, the diffusion model leverages the power of parallel computing to generate trajectories an order of magnitude faster (5c-5d).
V-D Real Robot Demonstration
We also test our method with the Robotarium [55] testbed to validate our method with real robot dynamics. Figure 6 shows the comparison between the RRT* heuristic and our method. We find the evader (red trajectory) equipped with our policy can reach the hideout and remain untracked by the searching team (blue and cyan trajectories) but the RRT* heuristic travels in a path with more cameras and is tracked by helicopter from $t\approx 90$. The result indicates that our data-driven learning based method can outperform the heuristics from the multi-agent interaction and map optimizations.
@@ V-D Real Robot Demonstration
(b) Diffusion-RL-Map (ours)
Figure 6: Pursuit-evasion trajectories: the evader, search party and helicopter trajectory are indicated with red, blue and cyan respectively. Evader is tracked from $t\approx 90$ with the RRT* heuristic but can escape from searching team with our method along the path to the hideout.
V-E Discussion
The significant disparity in goal-reaching rates between RL alone and our approaches indicates that RL struggles with long-horizon exploration in our domain, supporting our premise in §I that diffusion models can guide RL exploration. Additionally, Figure 3 reveals that SAC can become trapped in a local minimum. From this, we infer that our methods can achieve long-term planning by providing the RL agent with a diffusion-based global plan while retaining the agility to evade opponent agents in the short horizon. The cost map and time ablations in §V-B to §V-C demonstrate that our diffusion model can generate diverse samples and aid in downstream tasks such as cost map construction and path selection.

### v2

@@ IV-B RL Based Adversarial Behavior Learning
$\rho$ is the initial environment distribution for the start and hideout locations ($s_{0}\sim\rho$ is omitted from the derivation), $\pi^{d}$ is the diffusion global planner, and $\mathbf{w}$ is the generated waypoints to be tracked, $t_{g}$ is the number of waypoints. Equation (1) shows that we can decompose our adversarial evading learning problem into a sequential optimization problem, including a waypoint generator and a low-level waypoint tracker, both of which aim to maximize the discounted accumulative reward. However, the unconstrained random exploration of these policies will be subject to high variance from the partial observability; therefore, we confine our exploration in a more reasonable region described by the diffusion policy distribution in the previous section. We set reward terms to guide the evader to follow the waypoint and avoid detection in Equation (2) where $r_{g}$ is the reward obtained when reaching the waypoint, $r_{adv}$ is the detection penalty, and $r_{d}$ is the distance penalty to the next waypoint.
@@
TABLE I: Diffusion-RL benchmarks (mean $\pm$ std), G: global-view mode, L: local-view mode
Domain Prisoner Escape (G) Narco Interdiction (G) Prisoner Escape (L) Narco Interdiction (L)
Score $\uparrow$ Detection $\downarrow$ Goal-Reach $\uparrow$ Score $\uparrow$ Detection $\downarrow$ Goal-Reach $\uparrow$ Score $\uparrow$ Detection $\downarrow$ Goal-Reach $\uparrow$ Score $\uparrow$ Detection $\downarrow$ Goal-Reach $\uparrow$
@@
A-Star Heuristic 0.72$\pm$0.19 0.34$\pm$0.24 0.98$\pm$0.14 0.64$\pm$0.30 0.39$\pm$0.31 0.80$\pm$0.40 0.80$\pm$0.18 0.25$\pm$0.15 1.00$\pm$0.00 0.53$\pm$0.29 0.51$\pm$0.31 0.76$\pm$0.43
RRT-Star Heuristic 0.73$\pm$0.18 0.32$\pm$0.22 0.94$\pm$0.24 0.66$\pm$0.31 0.38$\pm$0.32 0.83$\pm$0.38 0.77$\pm$0.11 0.29$\pm$0.14 1.00$\pm$0.00 0.59$\pm$0.26 0.46$\pm$0.30 0.84$\pm$0.37
VO-Heuristic 0.74$\pm$0.11 0.32$\pm$0.14 0.98$\pm$0.14 0.64$\pm$0.10 0.43$\pm$0.12 1.00$\pm$0.00 0.73$\pm$0.12 0.33$\pm$0.15 0.99$\pm$0.10 0.64$\pm$0.10 0.43$\pm$0.11 1.00$\pm$0.00
@@
DDPG 0.23$\pm$0.34 0.73$\pm$0.40 0.07$\pm$0.26 0.31$\pm$0.32 0.65$\pm$0.37 0.00$\pm$0.00 0.23$\pm$0.34 0.73$\pm$0.40 0.07$\pm$0.26 0.26$\pm$0.36 0.70$\pm$0.42 0.00$\pm$0.00
SAC 0.78$\pm$0.12 0.06$\pm$0.11 0.13$\pm$0.34 0.49$\pm$0.35 0.43$\pm$0.41 0.01$\pm$0.10 0.73$\pm$0.12 0.10$\pm$0.15 0.05$\pm$0.22 0.51$\pm$0.40 0.41$\pm$0.47 0.00$\pm$0.00
Diffusion Only 0.78$\pm$0.09 0.28$\pm$0.12 1.00$\pm$0.00 0.78$\pm$0.13 0.26$\pm$0.15 1.00$\pm$0.00 0.79$\pm$0.10 0.26$\pm$0.13 1.00$\pm$0.00 0.78$\pm$0.13 0.26$\pm$0.15 1.00$\pm$0.00
@@
Diffusion-RL 0.90$\pm$0.09 0.12$\pm$0.10 0.96$\pm$0.20 0.89$\pm$0.11 0.12$\pm$0.10 0.94$\pm$0.24 0.90$\pm$0.12 0.11$\pm$0.10 0.90$\pm$0.30 0.89$\pm$0.11 0.11$\pm$0.11 0.91$\pm$0.29
Diffusion-RL-Map 0.94$\pm$0.08 0.07$\pm$0.08 0.98$\pm$0.14 0.94$\pm$0.07 0.07$\pm$0.08 1.00$\pm$0.00 0.95$\pm$0.09 0.05$\pm$0.07 0.95$\pm$0.22 0.94$\pm$0.08 0.06$\pm$0.07 0.95$\pm$0.22
@@ IV-C Cost Map Construction and Inference
Having discussed how the diffusion path planner could help the RL explore within the reward promising region in §IV-B and how the low-level SAC policy could be derived under the global path and initial environment distribution, we now take a step further and describe which global path sample should be drawn in each test case.
Even though the evasive agent does not have direct access to the camera or the terrain map information, this agent can still perform reverse spatial inference on the map by interacting with the learned policies to identify regions of heightened risk. Following this idea, we construct a cost map to reflect the risk of detection which we can then use to score the global paths.
Algorithm 4 describes our procedure. We run a certain number of episodes after the hierarchy is trained well (Line 4) and simultaneously record the evader location $\mathbf{X}$ and its distance to the closest searching agent $\mathbf{d}$ (Line 4). If the searching agent is within the evader’s field of view, we select out the evader locations $x_{i}$ with which the associated distance, $d_{i}$, is below a risk threshold $\epsilon$ and add a Gaussian distribution whose mean is $x_{i}$ and standard deviation $\sigma$ (Line 4-4). Next, we perform normalization and output the map.
Importantly, end-users can preview the global paths before execution or revise the cost map with the newly added global information, which adds more flexibility and interpretability compared with traditional fully MDP-based RL algorithms (Line 4). We can select the global path with the lowest cost based on the costmap (Line 4-4).
@@ 2       $\mathbf{x},\mathbf{d}=\mathbf{Env}(s_{0},\pi^{d},\pi^{l*})$
4             if $d_{i}<\epsilon$ see searching agent then
@@ V Results & Discussion
In this section, we will first benchmark our diffusion-RL framework for the global- and local-view evader against both heuristics and learning based baselines (see §V-A). Then we visualize our constructed costmap and show the interpretability, efficiency, and generalizability of our method in multiple ablation studies (see §V-BV-D). Next, we validate our hierarchy in the real robot testbed (see §V-E). Finally, we include a result discussion section (see §V-F).
Figure 3: Our Diffusion-RL trajectory vs SAC trajectory. Colorbar indicates the distance between the evader and closest pursuer and the purple line indicates the diffusion global path.
@@
Figure 4: Path Planning with the Costmap: The costmap (a) is constructed by correlating the agent’s risk of detection to its location on the map. We show in (b) that the agent can successfully identify where the cameras are. Given this costmap, the agent can select a path that best evades high-cost regions (c). Additional obstacles can be added ad-hoc (d) and a new path can be chosen (e). The grey areas indicate untraversable obstacles or danger zones.
V-A Achieving Long-Term Planning and Short-Term Evasion
We propose to use the following metrics to quantitatively evaluate our method and compare it against the baselines: detection, goal-reaching rate, and score ($\uparrow$ means higher is better and $\downarrow$ means lower is better in Table I). Detection is the average number of steps when the evader is detected for each episode. Goal-reaching rate is the number of episodes in which the evader reaches the goal within the time limit divided by the total number of episodes tested. The score metric is a weighted reward term that balances the two main objectives in our domains: reaching the goal ($\pm$50) and staying undetected by the opposing blue agents (-1 for each detection). Each metric is normalized by the min-max value of each category and we test the algorithms in 100 episodes with different random seeds from the training.
We compare our diffusion+RL method with learning based and non-learning-based baselines. A* and RRT* heuristics are from the open-sourced environment [12] where the evader follows the A* or RRT* path and shows some pre-scripted adversarial behaviors (e.g. heading to dense forest, changing velocity, etc.) when the search team is nearby. Velocity obstacle (VO) method is widely used in dynamic obstacle avoidance tasks which samples a velocity from the collision-free velocity space [47]. From Table I, we find these heuristics can reach the goal at a high rate in two view modes across both domains. However, the heuristic evader is detected by the search team much more than our methods, which leads to a lower score. The main reason is that the evader will be exposed to the cameras hidden from it, which may attract the dynamic search team and the hand-crafted evasive heuristic is not optimal for escaping.
Although RL-only approaches (DDPG [48] & SAC [49]) directly optimize the score as the reward function, our methods outperform them on average by 108% on the score metric across both domains. The SAC evader can be detected less than our diffusion-RL algorithm by 10.26% in prisoner escape domain, which is interesting and a red herring. Figure 3 shows a red colorbar for the distance between the evader and the closest agent in the pursuit team, and the diamond purple line indicates the diffusion planned global path. From the trajectory, we can see the SAC algorithm will lead the evader to the edge of the map where there is no camera but not the goal hideout. Instead, our method can loosely follow the diffusion waypoints (purple diamond) and simultaneously deviate a little to avoid the search team when they are close (indicated by magnet and orange dashed box). In addition, our method learns to reduce the evader speed when evading, which lowers the probability of detection, but maintains a higher speed when safer. We also ablate our diffusion-RL with the diffusion-only baseline (Diffuser [36]) and the 56.6% detection rate decreasing shows the low-level RL policy is crucial to adversarial evading behaviors.
V-B Interpretability and Flexibility
As discussed in §IV-C, we build a costmap based on the distance between evader and pursuer which implies detection risks. The costmap serves as an interpretable and memory-efficient way to allow our hierarchical agent to plan based on past experiences and will improve the flexibility in the inference stage compared with RL-only motion planning.
@@ V-B Interpretability and Flexibility
Figure 4a shows the costmap, where the high-intensity region indicates the evader is close to the search team which may lead to detection. We also find that the potential detection usually happens around the cameras (see Figure 4b), which is reasonable since the searching team will search the region after the evader is detected by the camera. Therefore, Figure 4c shows the evasive agent can select the best global path which receives lowest cost and can best conceal it from the camera array. At the inference stage, we can adjust the costmap with the new information we received. For example, we can add an undesired danger zone for the evader on the costmap (see Figure 4d) such that it will select another way to reach another hideout which will avoid the danger zone (see Figure 4e). We ablate our model with and without the costmap and show that the detection from map-based path selection is further decreased by half.
V-C Efficiency and Diversity in Guidance
@@
Figure 5a-5b compares the diffusion global paths trained on the datasets from RRT* and A*. We see the diffusion model encodes a more diverse distribution of the paths leading to the final hideouts. As a result, A* guided RL method can only receive a score of 0.78$\pm$0.09, which is worse than the performance of our diffusion-RL method (0.90$\pm$0.09). Table 5c shows the planning time in the inference stage for both the diffusion and RRT* algorithms. The diffusion model takes 85.7% less time to generate one trajectory compared with RRT*, which requires a map search. Furthermore, the difference is enlarged when planning more paths since we can draw samples from the diffusion model in parallel. In addition, Figure 5d shows that the diffusion-guided RL training uses less than a half time to reach the same reward level (save 25.2 hours), which greatly improves training time efficiency crucial to parameter tuning. We generate the results using a machine with 11th Gen Intel Core i9 processor and GeForce GTX 1660 Ti graphic card.
(a) A* Trajectories
(b) Diffusion Trajectories
@@
(c) Inference Time (mean±$\pm$±std)
(d) Training Time (Hours)
Figure 5: The paths from the diffusion model trained on RRT* are more diverse than A* paths (5a-5b). Compared to the traditional RRT* planner, the diffusion model leverages the power of parallel computing to generate trajectories an order of magnitude faster (5c). The diffusion-guided RL training also significantly decreases the training time (5d).
V-D Generailizability to New Camera Map
The high-level diffusion and low-level RL policies trained on one map can be generalized to other maps with different camera distributions with the only need to re-run costmap reconstruction process rather than the whole training pipeline (see Figure 6). We benchmark with the new camera map, and our results are shown in the Table II. Our method still outperforms the best baseline by 18.6% and outperforms 43.99% across all baselines in average, which shows our algorithm can be generalized to the new camera map without any retraining of diffusion or RL policies.
(a) New Camera Map
(b) New Costmap
Figure 6: (6a) - Cameras are highlighted. We use a different camera distribution in the test scenario (indicated by orange) from the one in the training scenario (indicated by blue). (6b) - we derive a new costmap reflecting the testing scenario camera distribution by running Algorithm 4.
TABLE II: Generalizability Benchmarks (mean $\pm$ std)
Domain (global-view) Prisoner Escape
Score $\uparrow$ Detection $\downarrow$ Goal-Reach. $\uparrow$
Non-Learning Approaches
A-Star Heuristic 0.694$\pm$0.185 0.376$\pm$0.229 0.970$\pm$0.171
RRT-Star Heuristic 0.701$\pm$0.179 0.367$\pm$0.220 0.970$\pm$0.171
VO-Heuristic 0.728$\pm$0.104 0.336$\pm$0.131 0.980$\pm$0.140
Learning Approaches
DDPG 0.136$\pm$0.241 0.847$\pm$0.264 0.070$\pm$0.255
SAC 0.670$\pm$0.203 0.195$\pm$0.226 0.130$\pm$0.336
Diffusion Only 0.767$\pm$0.080 0.292$\pm$0.100 1.000$\pm$0.000
Our Approaches
Diffusion-RL [ours] 0.864$\pm$0.111 0.155$\pm$0.116 0.940$\pm$0.237
Diffusion-RL-Map [ours] 0.910$\pm$0.075 0.113$\pm$0.093 1.000$\pm$0.000
V-E Real Robot Demonstration
We also test our method with the Robotarium [50] testbed to validate our method with real robot dynamics. Figure 7 shows the comparison among the RRT* heuristic, SAC, and our method. We find the evader (red trajectory) equipped with our policy can reach the hideout and remain untracked by the searching team (blue and cyan trajectories). In comparison, the SAC policy does not lead the evader to the hideout, and the RRT* heuristic travels a path near cameras and is tracked by the helicopter starting at $t\approx 90$. This validates our benchmarking results and indicates that our hierarchical data-driven learning-based method can outperform the baselines from the efficient diffusion-guided exploration, RL-based trial-and-error multi-agent interaction, and map optimizations.
@@ V-E Real Robot Demonstration
(b) SAC
(c) Diffusion-RL-Map
Figure 7: Pursuit-evasion trajectories: the evader, search party, and helicopter trajectory are indicated with red, blue, and cyan, respectively. We use a diffusion model to guide the evader to the hideout and explore escaping behaviors with RL, which performs better than RRT* heuristic with insufficient evasive behaviors and SAC method converging to a suboptimal place.
V-F Discussion
The significant disparity in goal-reaching rates between standalone RL and our approaches indicates that RL struggles with long-horizon exploration in large partially observable domains, supporting our premise in §I that diffusion models can guide RL exploration. Additionally, Figure 3 reveals that SAC can become trapped in a local minimum. From this, we infer that our methods can achieve long-term planning by providing the RL agent with a diffusion-based global plan while retaining the agility to evade opponent agents in the short horizon. The ablation studies in §V-B to §V-D demonstrate that our hierarchy can generate more interpretable paths, have more flexible motion plans and generalize to different test maps. Importantly, it significantly improves training and inference efficiency, which unlocks the potential on more applications in real-time tasks [51, 52].